OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. 18° 36° 54° 58°

OpenStudy (mathstudent55):

Angle: x Complement: 90 - x

OpenStudy (mathstudent55):

Sum of the complements: x + 90 - x = 90

OpenStudy (anonymous):

2x??

OpenStudy (mathstudent55):

Difference of their measures: x - (90 - x)

OpenStudy (mathstudent55):

90 exceeds x - (90 - x) by 72 That means: 90 - [x - (90 - x)] = 72

OpenStudy (mathstudent55):

Solve for x.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

i distribute the negitive?

OpenStudy (mathstudent55):

If x is more than 45, then subtract x from 90

OpenStudy (anonymous):

so it would be a

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

90 - [x - (90 - x)] = 72 90 - [x - 90 + x] = 72 90 - [2x - 90] = 72 90 - 2x + 90 = 72 -2x + 180 = 72 -2x = -108 x = 54 The measure of one of the angles is 54. Since complementary angles add to 90, the other angle is 90 - x. 90 - 54 = 36 Since they ask fot eh measure of the smaller angle, the answer is 36.

OpenStudy (anonymous):

thank you man that much simpler

OpenStudy (mathstudent55):

^the last sentence above should read: Since they ask for the measure of the smaller angle, the answer is 36.

OpenStudy (mathstudent55):

You're welcome.